Serena's father is twenty-six years older than serena. In ten years, the sum of the ages will be eighty. What are their present ages? Please Help Thank you.
Found 2 solutions by algebrahouse.com, richwmiller:
Answer by algebrahouse.com(1659)

"Serena's father is twenty-six years older than serena. In ten years, the sum of the ages will be eighty. What are their present ages?"

x = Serena's age now
x + 26 = father's age now

x + 10 = Serena's age in 10 years
x + 36 = father's age in 10 years {added 10 on to each one}

x + 10 + x + 36 = 80 {in ten years the sum of the ages will be 80}
2x + 46 = 80 {combined like terms}
2x = 34 {subtracted 46 from both sides}
x = 17 {divided both sides by 2}
x + 26 = 43 {substituted 17, in for x, into x + 26}

Serena is 17
her father is 43

For more help from me, visit: www.algebrahouse.com



Answer by richwmiller(17219)

s+26+10+s+10=80
2s+26=60
s+13=30
s=17
s+26=43